OpenStudy (anonymous):

evauluate the limit as x->infinity x+(x^2+3x-1)^/4x+5

OpenStudy (atlas):

You wrote (x^2 +3x-1)^ but you did not write any number after ^........are you missing something?

OpenStudy (anonymous):

Actually the whole polynom is ^2 sorry forgot the 2*

OpenStudy (anonymous):

evauluate the limit as x->infinity x+(x^2+3x-1)^2/4x+5 it goes like this

OpenStudy (atlas):

Expand the numerator alone and tell me.......Ask if you have any problem doing that

OpenStudy (anonymous):

actually i wanna have the solution of this kind of limit , not just the short answer i mean step by step, i can try to draw it if u have troubles understand the limit ..

OpenStudy (atlas):

I understand the question........i am trying you to work out the answer on your own

OpenStudy (anonymous):



OpenStudy (anonymous):

actually i do have an idea

OpenStudy (atlas):

oh this is different frm what you wrote b4

OpenStudy (anonymous):

by factorising the x inside the Decreased Square

OpenStudy (atlas):

Anyway........what's your idea?

OpenStudy (anonymous):

yeah this one is the limit , the drawn one , because i didnt knew how to write it

OpenStudy (atlas):

k.......whenever we have the questions of this sort involving numerator and denominator........we first find the highest power of x in the question and divide both the numerator and denominator by that

OpenStudy (anonymous):

Factorising the x^2 inside the Decreased Square both 3x and -1 goes 0 and we get 2x/4x+5 , then 5 goes 0 by dividing 5 with x and we get 2x/4x = 1/2

OpenStudy (atlas):

what is the highest power of x here

OpenStudy (anonymous):

2?

OpenStudy (atlas):

Nope x^2 is under square root sign...........So highest power is 1

OpenStudy (atlas):

Let us divide both numerator and denominator by x.....

OpenStudy (anonymous):

so do we have to factorise the highest power of x? right? to get it out of root square?

OpenStudy (anonymous):

factorise inside root square?

OpenStudy (atlas):

Nope factorization is not needed here

OpenStudy (anonymous):

How can we devide nomerator? Because there is x+(....) shouldnt we simplify it first? idk how this works

OpenStudy (atlas):

\[\frac{ x + \sqrt{(x^2 + 3x -1)}\ }{ 4x +5 }\] \[\frac{ x/x + \sqrt{(x^2/x^2 + 3x /x^2-1/x^2)}\ }{ 4x/x+5 }\]

OpenStudy (atlas):

Is this the question or .......is it x +( num/den)

OpenStudy (anonymous):

no the question is right as u wrote it

OpenStudy (anonymous):

The answer should be 1/2 final answer : lim x->infinity = 1/2 but i should only know the way

OpenStudy (anonymous):

are you sure we are able to divide x/x+(x2/x2+3x/x2−1/x2)/ 4x/x+5

OpenStudy (anonymous):

it goes 1+0 / 4x+5

OpenStudy (atlas):

It gets you (1 + 0 )/4 = 1/4

OpenStudy (anonymous):

yea it must be 1/2 the book i found this limit got the right answer and its 1/2 the short answer

OpenStudy (anonymous):

Maybe we need to factorize inside

OpenStudy (anonymous):

to get 2+0 /4 = 1/2

OpenStudy (atlas):

Sry.........it gives you (1+1)/4 = 2/4

OpenStudy (atlas):

check under the square root you get : (1+0+0) when you apply x->infinity

OpenStudy (anonymous):

Oh yes i see

OpenStudy (anonymous):

x2/x2 = 1

OpenStudy (atlas):

x/x =1 x^2/x^2 =1 3x/x^2 = 3/x = 0 (when x->infinity) 1/x^2 = 0 (when x->infinity) 4x/x =4 and 5/x^2 =0 (when x->infinity)

OpenStudy (atlas):

so you get ( 1+1 )/4 = 1/2

OpenStudy (anonymous):

yep thanks , one more question why 5 get divided by x^2? and not by x?

OpenStudy (atlas):

sry.......my mistake 5 gets divided by x and NOT x^2

OpenStudy (anonymous):

yeap no problems thanks alot for this quite helpfull answer

OpenStudy (atlas):

you are welcome